A population is estimated to have a standard deviation of 11. We want to estimate the population mean within 3, with a 90% level of confidence. (Use z Distribution Table.)

How large a sample is required? (Round the z-values to 2 decimal places. Round up your answer to the next whole number.)

Sample required is ____?

Answer by rothauserc(4718)

We use the following formula to solve this problem
:
E = z(a/2) * (s/square root(n)), where
:
E is the required difference between the sample mean and the population mean
:
z(a/2) is the positive z value for a(alpha)/2
:
s is the standard deviation of the population
:
n is the number for the sample size
:
n = ( z(a/2) * s / E )^2
:
a = 1 - (90/100) = 0.10
:
z(0.05) = 1.65
:
n = ( 1.65 * 11 / 3)^2 = 36.6025
:
*************************
the sample required is 37
*************************
: